Pacheo Corporation, which has only one product, has provided the following data concerning its most recent month of operations:

Selling price $55

Units in Beginning Inventory 500

Units Produced 3,100

Units Sold 3,300

Units in ending inventory 300

Variable Costs per unit

Direct Materials $19

Direct Labor $26

Variable manufacturing overhead $1

Variable Selling and Administrative expense $4

Fixed Costs:

Fixed Manufacturing Overhead $12,400

Fixed Selling and Administrative Costs $3,300

The company produces the same number of units every month, although the sales in units vary from month to month. The company 's variable costs per unit and total fixed costs have been constant from month to month.

Required:

a. What is the unit product cost for the month under variable costing? Show all work!

b. Prepare a contribution format income statement for the month using variable costing. Show all work!

c. Without preparing an income statement, determine the absorption costing net operating income for the month. (Hint: Use the reconciliation method)
Answer #1

Answer a.

Product cost under variable costing $            46

Calculated as

Calculation of per unit Variable Cost of Goods sold
Direct Material Per unit $            19
Direct Labor Per unit $            26
Variable Manufacturing overhead $               1
Variable Cost of Goods sold Per unit $            46

Answer b.

Variable Costing Income Statement
Sales $ 181,500
Variable Expenses
Variable Cost of Goods Sold = 3300*46 $ 151,800
Variable Selling and Administrative Expense = 3300*4= $    13,200
. $ 165,000
Contribution Margin = Sale-Variable expense = $    16,500
Fixed Expenses
Fixed manufacturing overhead $    12,400
Fixed selling and administrative expense $      3,300
Total Fixed Expenses $    15,700
Net operating income $          800

Answer c.

Reconciliation of variable costing and Absorption costing net operating incomes
Variable costing net operating income (loss) $          800
Less: Fixed manufacturing overhead costs deferred in inventory under absorption costing 200
Absorption costing net operating income (loss) $          600